ok, i'm stuck. I am trying to solve this quadratic equation by factoring. Can you help me?
The problem is:
x^2-11x+18=0
to find -11, I got -9 x -2 which is equal to =18
so I end up with (x-9)(x-2)=0 now where do i go from here?
is it x-9=0 x-2=0
+9 +9 +2 +2
so the answer is 2,9. but how do i check it.

Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

You've done fine. The solutions are 2,9
Put them back in the equation.
X^2-11x+18=0
X=2
4 - 22 +18=0. Yes
X=9
81 - 99 + 18=0. Yes
Done.

Answer by stanbon(75887)

ok, i'm stuck. I am trying to solve this quadratic equation by factoring. Can you help me?
The problem is:
x^2-11x+18=0
to find -11, I got -9 x -2 which is equal to =18
so I end up with (x-9)(x-2)=0 now where do i go from here?
is it x-9=0 x-2=0
+9 +9 +2 +2
so the answer is 2,9.
-----
That's right. You're doing just find.
--------------------
but how do i check it?
--------
Substitute each of the "answers" into
x^2-11x+18=0
Checking:: x = 2
2^2-11*2+18 = 0
4-22+18 = 0
0 = 0
--------------
Checking:: x = 9
9^2-11*9+18 = 0
81-99+18 = 0
-18+18 = 0
0 = 0
-------------
Cheers,
Stan H.
--------------